# QMeCha: quantum Monte Carlo package for fermions in embedding environments

### v1

@@ III.1.2 The antisymmetrized geminal power
To recall the relationship between the Pfaffian wavefunction and antisymmetrized geminal power (AGP) wavefunction[95, *col+65jmp, 97, 98] let us consider a closed shell system with an even number of electrons. By removing in the P matrix defined in eq. 20, the coupling between the electrons with parallel spin ($\mathcal{Z}^{\uparrow}=\mathcal{Z}^{\downarrow}=0$), the Pfaffian reduces to the determinant of the geminal matrix G[88, 89],
@@
which corresponds to the AGP wavefunction[95, *col+65jmp, 97, 98] for closed shell systems. The elements of the geminal matrix are thus those defined in eq. 23 and describe the coupling between electrons of opposite spin in a pure singlet state.
@@ III.2 Electron-Positron Correlations
The limitation of this type of factorization lies in the fact that, while atomic basis sets are a good basis for electrons that form bound states with the nuclei, they are not an appropriete representation for positrons that are repelled by the atomic nuclei and attracted by the electrons with which they can form metastable states before annihilation[48, 50, 49, 104, 105, 106, 107, 108, 87, 109, 110].
Thus, for the case of many electrons and one positron, an alternative way to construct the positronic wavefunction is to assume that while the electronic wavefunction $\psi_{e}(\bar{\textbf{x}}^{e})$ describes the spin and angular symmetries of the electrons in the field of the nuclei, $\psi_{p}(\textbf{x}^{p},\bar{\textbf{x}}^{e})$ depends only on the distances between the single positron and electrons, and not on the nuclear coordinates; with the constraint of being symmetric for the exchange of any electronic coordinate.
One way to achieve this is by constructing $\psi_{p}(\textbf{x}^{p},\bar{\textbf{x}}^{e})$ through the ‘Positronium’ basis set[82, 52, 38], which explicitly depends on the electron-positron distance. As a matter of fact, it can be easily shown that the ground state of a system of one electron and one positron, i.e. the Positronium (Ps), can be exactly described by an exponential function of the electron-positron distance $r^{ep}=|\textbf{x}^{e}-\textbf{x}^{p}|$:
$\varphi\left(\textbf{x}^{ep}\right)=r^{ep}R\left(r^{ep}\right)Y_{l}^{m}(\theta^{ep},\phi^{ep}),$ (31)
@@
$\psi_{p}(\textbf{x}^{p},\bar{\textbf{x}}^{e})=\prod_{i=1}^{N_{e}}\phi_{p}(\textbf{r}^{ep}_{i}),$ (32)
@@ III.3 Jastrow factors
$\mathcal{J}(\bar{\textbf{x}})=\text{exp}\{J_{c}(\bar{\textbf{x}})+J_{d}(\bar{\textbf{r}})\}$ (34)
both depending on the fermionic and nuclear degrees of freedom (that have been omitted in the equation), of which the first contains an explicit description of the wavefunctions’ cusp conditions (both for fermion-nucleus and fermion-fermion), while the second is used to describe the few-body correlation effects between the fermions in the field of the nuclei. This Jastrow is a generalization[38, 39] of that introduced in refs. 98, 103.
@@ III.3.1 General cusp functions
The Jastrow $J_{c}(\bar{\textbf{x}})$ term that is used to describe the many particles’ cusp conditions is written as the linear combination
@@ III.3.2 Dynamical Jastrow factor
As anticipated at the beginning of section III.3 the dynamical part $J_{d}(\bar{\textbf{r}})$ of the Jastrow factor is a generalization of that introduced in refs. 98, 103, based on an expansion in $Q$ non-normalized atomic orbitals $\chi_{\nu}(\textbf{r})$ (the atomic index is included in the orbital index $\mu$ or $\nu$ and will always be omitted in the following sections).[38, 39]
@@ III.4 Wavefunction of drudons
The QDO wavefunction $\Psi_{Q}(\bar{\textbf{r}}^{d};\bar{\textbf{R}}^{c})$ is constructed to depend only on the coordinates of the Drudons and parametrically on the centers of QDOs.
@@
An alternative approach to the QDO wavefunction, introduced in ref. 41 and is built as the Hartree product of
@@
$\phi_{i}(\textbf{r}^{d}_{i})=\sum_{q=1}^{Q}\alpha^{i}_{q}\varphi_{q}(\textbf{r}_{i})$ (43)
@@ III.6 Basis sets
the radial part of the wavefunction can be written as a linear combination of $p$ different primitive functions.
@@
where the primitive functions are in general written as
$P_{n}(r_{a})=\mathcal{N}_{n,l}(\zeta)r_{a}^{n}e^{\alpha(\zeta,r_{a})}.$ (48)
@@ IV Quantum Monte Carlo methods
The quantum Monte Carlo (QMC) methods[28, 29] implemented in QMeCha are variational Monte Carlo (VMC) and diffusion Monte Carlo (DMC). Within the VMC framework, it is possible to do a full optimization of the variational space using the Stochastic Reconfiguration (SR) approach introduced by S. Sorella in ref. 118, 29.
@@ IV.1 Variational Monte Carlo
It is important to notice that, from the definition of the variance in eq. 56, all QMC methods satisfy the zero variance principle[29], ie. if the trial wavefunction $\Psi_{T}(\bar{\textbf{r}};\boldsymbol{\alpha})$ is an eigenfunction of the Hamiltonian[126, 127], the local energy function $\text{E}_{l}(\bar{\textbf{r}};\boldsymbol{\alpha})$ becomes a constant that corresponds to the eigenvalue associated to $\Psi_{T}(\bar{\textbf{r}};\boldsymbol{\alpha})$, and thus the variance is null.
@@
The parameters $\delta_{e}$, $\delta_{p}$, and $\delta_{d}$ are amplitudes used respectively for the electrons, positrons, and drudons, while $m_{e}$, $m_{p}$ and $\mu_{i}$ are the corresponding masses. Within the code, the $\delta_{e}$, $\delta_{p}$, and $\delta_{d}$ parameters are optimized during the thermalization process by converging the acceptance probability of the MC moves to the value of 50%, which is the rule of thumb, that has the purpose of balancing the acceptance rate of the single particle moves and the correlation between configurations and thus between local observable evaluations. In order to further reduce the correlation, the two-step algorithm developed in ref. 132 is also applied to split the acceptance probability between the determinant part of the total wavefunction plus the one-body Jastrow, and the two- the three- and four body Jastrow factors implemented in the code[97, 38, 40, 41] (See Section III.3).
@@ IV.2 Optimization methods
can be computed via a conjugate gradient (CG) algorithm with parallel and distributed, nearly independent, computational tasks. Since statistical noise can deteriorate the stability of the SR optimization, due to the fact that the eigenvalues of the covariance matrices can be very small, amplifying the noise in the stochastic forces, here we apply a constant positive shift of the diagonal matrix elements to regularize the matrix, as previously suggested in refs. 29, 131, so that $\textbf{S}^{\prime}=\textbf{S}+\epsilon\textbf{I}$ with $\epsilon\in[0.0001,0.01]$ depending on the different scales of the parameters[29, 131]. Furthermore, in order to improve the convergence of the CG method, the S matrices are preconditioned according to the procedure reported in ref. 29. This optimization procedure is also implemented within the framework of the correlated sampling technique. In this case, the code measures the overlap between the wavefunction used for the initial sampling and the wavefunction at a given step and decides based on a threshold when to automatically resample if needed.
@@ IV.3 Diffusion Monte Carlo
In QMeCha we have implemented a fixed-node diffusion Monte Carlo algorithm inspired by the one first published by Umrigar et al. in ref. 119.
@@
The fixed node procedure in QMeCha differs from that in ref. 119 by a set of more modern features that have been introduced to improve the estimation of energy difference with pseudopotential calculations. In particular, to integrate the pseudopotential component determinant locality approximation (DLA)[123] is used to increase stability and reproducibility of the DMC energies independently on the type of Jastrow[36]. Second, following the work of Anderson et al. [121, 120] in order to reduce the size-consistency error of the DMC calculations[122], a single particle energy cut-off algorithm has been introduced, removing the original dependency of the branching factor on the effective time step[120] due to the fact that this increases the size-consistency error.
The efficiency of the algorithm is shown in the code comparison article ref. 36.
@@ V Code Structure, Computational details and Efficiency
The QMeCha code is a modular tool written in Fortran 2008 with two levels of parallelization achieved with both Message Passing Interface (MPI)[135] and shared-memory Open Multi-Processing (OpenMP)[136]. The general structure of the code is briefly represented in Fig. 3.
The code is available under Creative Commons Attribution-NonCommercial-NoDerivatives (CC BY-NC-ND) license through the GitHub repository github.com/QMeCha.
The GitHub repository includes the QMeCha a repository of tools for the initialization of the wavefunction files github.com/QMeCha/QMeCha_tools and the open version of the code github.com/QMeCha/QMeCha_code.
The first step is the initialization of the MPI/OpenMP environment. Afterwards, the code takes as input a series of arguments that are used to specify the input files that include the systems’ properties, such as charge, spin, and atomic coordinates, and the basis set files. The input files for the fermionic system and the embedding environment are separated. These input files can also be specified in the main input file that includes the details of the QMC runs that have to be executed – that is indicated as \${inputfile} in Fig. 3 – overwriting the information from the input arguments.
@@
Within the code, after the initializations of the molecular and embedding systems and of the basis sets and wavefunctions, the code calls the module that executes the VMC, DMC or wavefunction optimization. Each molecular system property, each QMC method, and each wavefunction component are handled by separate Fortran modules progressively linked together in order to guarantee a simple way to progressively expand each main element of the code, without conflicts.
@@
Figure 4: Weak scaling tests for the system of 30 water molecules all described as electronic systems. The original geometry is taken from ref. 60 (sites.uw.edu/wdbase) and is shown in Fig. 1a.
@@ VI Applications
Figure 5: a. Difference between the binding energies (BE) of the dimers computed with FN-DMC and DFT (PBE0+MBD) and the LNO-CCSD(T) calculations. For the LNO-CCSD(T) calculations we also report the estimated uncertainty that includes the basis set extrapolation and the error due to the orbital localization (see ref. 43 for further details), indicated with the different $\sigma,2\sigma,3\sigma$ curves. These are only used as a guide for the eye to underline the dimers for which there is the largest discrepancy between FN-DMC and LNO-CCSD(T). b. Relative error (in percentage) of the FN-DMC and DFT (PBE0+MBD) calculations defined with respect to the LNO-CCSD(T) ones as $|(\text{BE}-\text{BE}_{\text{LNO-CCSD(T)}})/\text{BE}_{\text{LNO-CCSD(T)}}|$. In this case the errors on the FN-DMC and PBE0+MBD calculations are obtained through the propagation of error that also includes the uncertainty in the LNO-CCSD(T) results. On the left we report the average relative absolute errors for FN-DMC and PBE0+MBD respectively in blue and red. In panels c, d, e, f we display the structures of the conformers F1I2, L2I2, SF1I2 and SF3I3 respectively, that show the larges discrepancies between FN-DMC and LNO-CCSD(T) results, as shown in panel a. Data is taken from refs. 43, 138.
@@ VI.1 Van der Waals interactions in macro-molecules
The two reference methods that are usually used to provide accurate estimates for small and medium trial systems[139, 141, 142, 143, 144, 145, 146, 147, 148] are FN-DMC[149, 29] and CCSD(T)[20]. On one hand, CCSD(T)[20] is regarded as the ‘gold standard’ method of quantum Chemistry able to converge the dynamical correlation for energy and force estimations for a large variety of chemical compounds. On the other hand, also FN-DMC is expected to give similar or even more accurate results, since the method is able to recover all the dynamical electronic correlation responsible for such bindings. In fact, during the diffusion process the wavefunction amplitudes’ responsible for the bonding can converge to the exact result, while the nodal structure of two vdW-interacting fragments does not vary significantly with respect to the non-interacting case.
@@
Clearly, some of this discrepancy could stem from the inconsistency in the description of the nodal structures of the vdW-interacting fragments with those in the non-interacting limit that, in some cases, such as the water dimer[154, 155], require multi-determinantal approaches or backflow transformations to converge. However, as also shown in a recent work by Shi et al. [148], for the dispersive interacting systems of the S66[143, 144] dataset the FN-DMC results have only a mean average deviation of 0.09 kcal/mol with respect to the CCSD(cT)[153] calculations, that differ from CCSD(T) only on the improved approximation to the triple particle-hole excitation amplitudes. Furthermore, within these energy differences the convergence of CC with respect to the order of explicit and perturbative excitations included has been shown to not converge monotonically[156]. Therefore, for these small interaction energies that become crucial in progressively large molecular systems, the puzzle of the differences between FN-DMC and CCSD(T) still remains unsolved, calling for the improvement of the state-of-the-art of both CC and FN-DMC methods.
To shed some light on this puzzle, recently QMeCha was used to obtain FN-DMC reference calculations, compared to LNO-CCSD(T)[44, 45, 46, 47], for a set of molecular dimers each containing hundreds of electrons (see Fig. 5) in ref. 43.
@@
The interactions between the chosen subset of the AQM molecules and the two small monomers have the property of covering a large range of different and multiple vdW interaction types within each dimer, such as $\pi$-$\pi$ stacking, hydrogen bonding, electrostatics, and dispersion interactions.
@@
In Fig. 5a, we have summarized the resulting differences between the binding energies (BE) computed with FN-DMC and PBE0+MBD with respect to the LNO-CCSD(T) results. The conformers in the QUID dataset are originally catalogued through a string in which the first two letters indicate the main structural property of the macro-molecule, i.e. folded (F), linear (L), semi-folded (SF), the first number indicates the number of the conformer, the letters I and B indicate the interaction with imidazole and benzene respectively and the last number indicates the binding site. Further details can be found in ref. 43. In this panel, we also display the uncertainty on the LNO-CCSD(T) results that come from the localization and basis set extrapolation errors. In Fig. 5b, we report the relative error $\left|(\text{BE}-\text{BE}_{\text{LNO-CCSD(T)}})/\text{BE}_{\text{LNO-CCSD(T)}}\right|$ in percentage on the binding energies. Here, the estimated uncertainties on the LNO-CCSD(T) extrapolated results is propagated on both the FN-DMC and the PBE0+MBD reported errors.
@@
From these results, it is evident that the biggest discrepancies can be found in those dimers involving the macromolecule interacting with imidazole and presenting hydrogen bonds. In particular, in Fig. 5a we highlight four dimers that are displayed in Figs. 5c,5d,5e,5f.
@@
In all these conformers, there appears to be a hydrogen bond between the nitrogen atom of one monomer and the hydrogen of the other, sometimes the donor is imidazole (Figs. 5d,5f) and sometimes the macro-molecule (Figs. 5c,5e). Clearly, it is impossible to exclude that this effect comes from the inconsistency between the description of the nodal structure of the bonded dimer and of the non-interacting one, thus the question of how the convergence of the nodal surface affects the accuracy of the descriptions of such interactions still remains an open question that deserves further investigation.
@@
In general, these results confirm what was also observed in refs. 146, 147, 148 that FN-DMC with a Slater determinant wavefunction tends to overestimate the binding energies of hydrogen-bonded dimers by about 0.1 kcal/mol, while the results obtained for dispersive interacting systems can achieve significant accuracy, and can clearly be used as reference calculation for the construction of less expensive computational methods.
@@
Table 1: Non-relativistic total energies (in Hartree) of atoms interacting with a positron ($e^{+}$) or with positronium (Ps). In parenthesis, we report the symmetry state of the electrons. All calculations are done with the SD or AGP wavefunction and the electron-positron wavefunction in eq. 32. SP indicated the single-pairing function that corresponds to one anti-symmetrized explicitly correlated pairing function from ref. 158, while MP is the linear combination of SP functions. Full data in ref. 38.
@@
CI -0.78874(60)a -24.83056b -38.05362b -75.28127b -100.001817c
@@
aFCI extrapolation from ref. 85. bFCI limit with higher momentum corrections from ref. 163. cMRCI calculation from ref. 164.

### v2

@@ III.1.2 The antisymmetrized geminal power
To recall the relationship between the Pfaffian wavefunction and antisymmetrized geminal power (AGP) wavefunction[106, *col+65jmp, 108, 109] let us consider a closed shell system with an even number of electrons. By removing in the P matrix defined in eq. 21, the coupling between the electrons with parallel spin ($\mathcal{Z}^{\uparrow}=\mathcal{Z}^{\downarrow}=0$), the Pfaffian reduces to the determinant of the geminal matrix G[99, 100],
@@
which corresponds to the AGP wavefunction[106, *col+65jmp, 108, 109] for closed shell systems. The elements of the geminal matrix are thus those defined in eq. 24 and describe the coupling between electrons of opposite spin in a pure singlet state.
@@ III.2 Electron-Positron Correlations
The limitation of this type of factorization lies in the fact that, while atomic basis sets are a good basis for electrons that form bound states with the nuclei, they are not an appropriate representation for positrons that are repelled by the atomic nuclei and attracted by the electrons with which they can form metastable states before annihilation[59, 61, 60, 115, 116, 117, 118, 119, 98, 120, 121].
Thus, for the case of many electrons and one positron, an alternative way to construct the positronic wavefunction is to assume that while the electronic wavefunction $\psi_{e}(\bar{\textbf{r}}^{e})$ describes the spin and angular symmetries of the electrons in the field of the nuclei, $\psi_{p}(\textbf{r}^{p},\bar{\textbf{r}}^{e})$ depends only on the distances between the single positron and electrons, and not on the nuclear coordinates; with the constraint of being symmetric for the exchange of any electronic coordinate.
One way to achieve this is by constructing $\psi_{p}(\textbf{r}^{p},\bar{\textbf{r}}^{e})$ through the ‘Positronium’ basis set[93, 63, 47], which explicitly depends on the electron-positron distance. As a matter of fact, it can be easily shown that the ground state of a system of one electron and one positron, i.e. the Positronium (Ps), can be exactly described by an exponential function of the electron-positron distance $r^{ep}=|\textbf{r}^{e}-\textbf{r}^{p}|$:
$\varphi\left(\textbf{r}^{ep}\right)=r^{ep}R\left(r^{ep}\right)Y_{l}^{m}(\theta^{ep},\phi^{ep}),$ (32)
@@
$\psi_{p}(\textbf{r}^{p},\bar{\textbf{r}}^{e})=\prod_{i=1}^{N_{e}}\phi_{p}(\textbf{r}^{ep}_{i}),$ (33)
@@ III.3 Jastrow factors
$\mathcal{J}(\bar{\textbf{r}})=\text{exp}\{J_{c}(\bar{\textbf{r}})+J_{d}(\bar{\textbf{r}})\}$ (35)
both depending on the fermionic and nuclear degrees of freedom (that have been omitted in the equation), of which the first contains an explicit description of the wavefunctions’ cusp conditions (both for fermion-nucleus and fermion-fermion), while the second is used to describe the few-body correlation effects between the fermions in the field of the nuclei. This Jastrow is a generalization[47, 48] of that introduced in refs. 109, 114.
@@ III.3.1 General cusp functions
The Jastrow $J_{c}(\bar{\textbf{r}})$ term that is used to describe the many particles’ cusp conditions is written as the linear combination
@@ III.3.2 Dynamical Jastrow factor
As anticipated at the beginning of section III.3 the dynamical part $J_{d}(\bar{\textbf{r}})$ of the Jastrow factor is a generalization of that introduced in refs. 109, 114, based on an expansion in $Q$ non-normalized atomic orbitals $\chi_{\nu}(\textbf{r})$ (the atomic index is included in the orbital index $\mu$ or $\nu$ and will always be omitted in the following sections).[47, 48]
@@ III.4 Wavefunction of drudons
The QDO wavefunction $\Psi_{d}(\bar{\textbf{r}}^{d};\bar{\textbf{R}}^{c})$ is constructed to depend only on the coordinates of the Drudons and parametrically on the centers of QDOs.
@@
An alternative approach to the QDO wavefunction, introduced in ref. 50 and is built as the Hartree product of
@@
$\phi_{i}(\textbf{r}^{d}_{i})=\sum_{q=1}^{Q}\alpha^{i}_{q}\varphi_{q}(\textbf{r}^{d}_{i})$ (44)
@@ III.6 Basis sets
the radial part of the wavefunction is written as a linear combination of $p$ different primitive functions.
@@
written as
$P_{n}(r_{a})=\mathcal{N}_{n,l}(\zeta)r_{a}^{n}e^{\alpha(\zeta,r_{a})},$ (49)
@@ IV Quantum Monte Carlo methods
The quantum Monte Carlo (QMC) methods[28, 29] implemented in QMeCha are variational Monte Carlo (VMC) and diffusion Monte Carlo (DMC). Within the VMC framework, it is possible to do a full optimization of the variational space using the Stochastic Reconfiguration (SR) approach introduced by S. Sorella in ref. 129, 29.
@@ IV.1 Variational Monte Carlo
It is important to notice that, from the definition of the variance in eq. 57, the VMC method presented here satisfies the zero variance principle[29], ie. if the trial wavefunction $\Psi_{T}(\bar{\textbf{r}};\boldsymbol{\alpha})$ is an eigenfunction of the Hamiltonian[135, 136], the local energy function $\text{E}_{l}(\bar{\textbf{r}};\boldsymbol{\alpha})$ becomes a constant that corresponds to the eigenvalue associated to $\Psi_{T}(\bar{\textbf{r}};\boldsymbol{\alpha})$, and thus the variance is null.
@@
The parameters $\delta_{e}$, $\delta_{p}$, and $\delta_{d}$ are amplitudes used respectively for the electrons, positrons, and drudons, while $m_{e}$, $m_{p}$ and $\mu_{i}$ are the corresponding masses. Within the code, the $\delta_{e}$, $\delta_{p}$, and $\delta_{d}$ parameters are optimized during the thermalization process by converging the acceptance probability of the MC moves to the value of 50%, which is the rule of thumb, that has the purpose of balancing the acceptance rate of the single particle moves and the correlation between configurations and thus between local observable evaluations. In order to further reduce the correlation, the two-step algorithm developed in ref. 141 is also applied to split the acceptance probability between the determinant part of the total wavefunction plus the one-body Jastrow, and the two- the three- and four body Jastrow factors implemented in the code[108, 47, 49, 50] (See Section III.3).
@@ IV.2 Optimization methods
can be computed via a conjugate gradient (CG) algorithm with parallel and distributed, nearly independent, computational tasks. Since statistical noise can deteriorate the stability of the SR optimization, due to the fact that the eigenvalues of the covariance matrices can be very small, amplifying the noise in the stochastic forces, here we apply a constant positive shift of the diagonal matrix elements to regularize the matrix, as previously suggested in refs. 29, 140, so that $\textbf{S}^{\prime}=\textbf{S}+\epsilon\textbf{I}$ with $\epsilon\in[0.0001,0.01]$ depending on the different scales of the parameters[29, 140]. Furthermore, in order to improve the convergence of the CG method, the S matrices are preconditioned according to the procedure reported in ref. 29. This optimization procedure is also implemented within the framework of the correlated sampling technique. In this case, the code measures the overlap between the wavefunction used for the initial sampling and the wavefunction at a given step and decides based on a threshold when to automatically resample if needed.
@@ IV.3 Diffusion Monte Carlo
In QMeCha we have implemented a fixed-node diffusion Monte Carlo algorithm inspired by the one first published by Umrigar et al. in ref. 130.
@@
The fixed node procedure in QMeCha differs from that in ref. 130 by a set of more modern features that have been introduced to improve the estimation of energy difference with pseudopotential calculations. In particular, to integrate the pseudopotential component, determinant locality approximation (DLA)[132] is used to improve stability and reproducibility of the DMC energies independently on the type of Jastrow[36], with respect to locality approximation (LA)[144]. T-moves[31] and determinant T-moves[36], that are more stable but computationally more expensive approaches have still to be implemented in the code, together with Lattice Regularized diffusion Monte Carlo (LRDMC)[145, 31].
Second, following the work of Anderson et al. [52, 53] in order to reduce the size-consistency error of the DMC calculations[131], a single particle energy cut-off algorithm has been introduced, removing the original dependency of the branching factor on the effective time step[53] due to the fact that this increases the size-consistency error.
The efficiency of the algorithm is shown in a code comparison investigation presented in ref. 36.
@@ V Code Structure, Computational details and Efficiency
The QMeCha code is a modular tool written in Fortran 2008 with two levels of parallelization achieved with both Message Passing Interface (MPI)[146] and shared-memory Open Multi-Processing (OpenMP)[147].
The code is available under Creative Commons Attribution-NonCommercial-NoDerivatives (CC BY-NC-ND) license through the GitHub repository github.com/QMeCha. The GitHub repository of QMeCha includes a repository of tools for the initialization of the wavefunction files github.com/QMeCha/QMeCha_tools and the open version of the code github.com/QMeCha/QMeCha_code.
The general structure of the code is briefly represented in Fig. 3. The first step is the initialization of the MPI/OpenMP environment. Afterwards, the code takes as input a series of arguments that are used to specify the input files that include the systems’ properties, such as charge, spin, and atomic coordinates, and the basis set files. The input files for the fermionic system and the embedding environment are separated. These input files can also be specified in the main input file that includes the details of the QMC runs that have to be executed – that is indicated as \${inputfile} in Fig. 3 – overwriting the information from the input arguments. A simple example of the main input files of QMeCha can be found in Fig. 4 for the case of the Oxygen atom described with the ccECP pseudopotential[81, *tra+14jctc, *tra+15jcp, *kro+16prb, *tra+17jcp, *tra+05jcp2, *ben+17jcp, *ben+18jcp, *ann+18jcp, *bur+07jcp, *bur+08jcp].
Figure 4: Main input files of the QMeCha code. a. Molecular system file. b. pseudopotential file (using the ccECP pseudopotential for Oxygen[81, *tra+14jctc, *tra+15jcp, *kro+16prb, *tra+17jcp, *tra+05jcp2, *ben+17jcp, *ben+18jcp, *ann+18jcp, *bur+07jcp, *bur+08jcp]). c. Wave function specifications, that include the type of wavefunction (RSD), the type of one- and two- body cusp functions (eq. 39) and the $c$-pVDZ basis set (truncated) for Oxygen, and the atomic Jastrow basis set.
@@
Within the code, after the initializations of the molecular and embedding systems and of the basis sets and wavefunctions, the code calls the module that executes the VMC, DMC or wavefunction optimization procedures. Each molecular system property, each QMC method, and each wavefunction component are handled by separate Fortran modules progressively linked together in order to guarantee a simple way to progressively expand each main element of the code, without conflicts.
@@
Figure 5: Weak scaling tests for the system of 30 water molecules all described as electronic systems. The original geometry is taken from ref. 71 (sites.uw.edu/wdbase) and is shown in Fig. 1a.
@@ VI Applications
Figure 6: a. Difference between the binding energies (BE) of the dimers computed with FN-DMC and DFT (PBE0+MBD) and the LNO-CCSD(T) calculations. For the LNO-CCSD(T) calculations we also report the estimated uncertainty that includes the basis set extrapolation and the error due to the orbital localization (see ref. 54 for further details), indicated with the different $\sigma,2\sigma,3\sigma$ curves. These are only used as a guide for the eye to underline the dimers for which there is the largest discrepancy between FN-DMC and LNO-CCSD(T). b. Relative error (in percentage) of the FN-DMC and DFT (PBE0+MBD) calculations defined with respect to the LNO-CCSD(T) ones as $|(\text{BE}-\text{BE}_{\text{LNO-CCSD(T)}})/\text{BE}_{\text{LNO-CCSD(T)}}|$. In this case the errors on the FN-DMC and PBE0+MBD calculations are obtained through the propagation of error that also includes the uncertainty in the LNO-CCSD(T) results. On the left we report the average relative absolute errors for FN-DMC and PBE0+MBD respectively in blue and red. In panels c, d, e, f we display the structures of the conformers F1I2, L2I2, SF1I2 and SF3I3 respectively, that show the larges discrepancies between FN-DMC and LNO-CCSD(T) results, as shown in panel a. Data from M. Puleva, L. Medrano Sandonas, B. D. Lőrincz, J. Charry, D. M. Rogers, P. R. Nagy and A. Tkatchenko Nat. Communications 16, 8583 (2025). ; licensed under a Creative Commons Attribution ($C$ BY) license.[54, 149].
@@ VI.1 Van der Waals interactions in macro-molecules
The two reference methods that are usually used to provide accurate estimates for small and medium size systems[150, 152, 153, 154, 155, 156, 157, 158, 159] are FN-DMC[160, 29] and CCSD(T)[20]. On one hand, CCSD(T)[20] is regarded as the ‘gold standard’ method of quantum Chemistry able to converge the dynamical correlation for energy and force estimations for a large variety of chemical compounds. On the other hand, also FN-DMC is expected to give similar or even more accurate results, since the method is able to recover all the dynamical electronic correlation responsible for such bindings.
@@
Clearly, some of this discrepancy stems from the inconsistency in the description of the nodal structures of the vdW-interacting fragments with those in the non-interacting limit that, in some cases, such as the water dimer[165, 166], require multi-determinantal approaches or backflow transformations to converge. However, as also shown in a recent work by Shi et al. [159], for the dispersive interacting systems of the S66[154, 155] dataset the FN-DMC results have only a mean average deviation of 0.09 kcal/mol with respect to the CCSD(cT)[164] calculations, that differ from CCSD(T) only on the improved approximation to the triple particle-hole excitation amplitudes. Furthermore, within these energy differences the convergence of CC with respect to the order of explicit and perturbative excitations included has been shown to not converge monotonically[167]. Therefore, for these small interaction energies that become crucial in progressively large molecular systems, the puzzle of the differences between FN-DMC and CCSD(T) still remains unsolved, calling for the improvement of the state-of-the-art of both CC and FN-DMC methods.
To shed some light on this puzzle, recently QMeCha was used to obtain FN-DMC reference calculations, compared to LNO-CCSD(T)[55, 56, 57, 58], for a set of molecular dimers each containing hundreds of electrons (see Fig. 6) in ref. 54.
@@
The interactions between the chosen subset of the AQM molecules and the two small monomers have the property of including multiple vdW interaction types within each dimer, such as $\pi$-$\pi$ stacking, hydrogen bonding, electrostatics, and dispersion interactions.
@@
In Fig. 6a, we have summarized the resulting differences between the binding energies (BE) computed with FN-DMC and PBE0+MBD with respect to the LNO-CCSD(T) results. The conformers in the QUID dataset are originally catalogued through a string in which the first two letters indicate the main structural property of the macro-molecule, i.e. folded (F), linear (L), semi-folded (SF), the first number indicates the number of the conformer, the letters I and B indicate the interaction with imidazole and benzene respectively and the last number indicates the binding site. Further details can be found in ref. 54. In this panel, we also display the uncertainty on the LNO-CCSD(T) results that come from the localization and basis set extrapolation errors. In Fig. 6b, we report the relative error $\left|(\text{BE}-\text{BE}_{\text{LNO-CCSD(T)}})/\text{BE}_{\text{LNO-CCSD(T)}}\right|$ in percentage on the binding energies. Here, the estimated uncertainties on the LNO-CCSD(T) extrapolated results is propagated on both the FN-DMC and the PBE0+MBD reported errors.
@@
From these results, it is evident that the biggest discrepancies can be found in those dimers involving the macromolecule interacting with imidazole and presenting hydrogen bonds. In particular, in Fig. 6a we highlight four dimers that are displayed in Figs. 6c,6d,6e,6f.
@@
In all these conformers, there appears to be a hydrogen bond between the nitrogen atom of one monomer and the hydrogen of the other, sometimes the donor is imidazole (Figs. 6d,6f) and sometimes the macro-molecule (Figs. 6c,6e). Clearly, it is impossible to exclude that this effect comes from the inconsistency between the description of the nodal structure of the bonded dimer and of the non-interacting one, thus the question of how the convergence of the nodal surface affects the accuracy of the descriptions of such interactions still remains an open question that deserves further investigation.
@@
In general, these results confirm what was also observed in refs. 157, 158, 159 that FN-DMC with a Slater determinant wavefunction tends to overestimate the binding energies of hydrogen-bonded dimers by about 0.1 kcal/mol, while the results obtained for dispersive interacting systems can achieve significant accuracy, and can clearly be used as reference calculation for the construction of less expensive computational methods.
@@
Table 1: Non-relativistic total energies (in Hartree) of atoms interacting with a positron ($e^{+}$) or with positronium (Ps). In parenthesis, we report the symmetry state of the electrons. All calculations are done with the SD or AGP wavefunction and the electron-positron wavefunction in eq. 33. SP indicated the single-pairing function that corresponds to one anti-symmetrized explicitly correlated pairing function from ref. 169, while MP is the linear combination of SP functions. Data partially taken from J. A. Charry Martinez, M. Barborini and A. Tkatchenko J. Chem. Theory Comput. 18, 4, 2267–2280 (2022); licensed under a Creative Commons Attribution ($C$ BY) license.[47].
@@
CI -0.78874(60) [^1] FCI extrapolation from ref. 96. -24.83056222FCI limit with higher momentum corrections from ref. 174. -38.05362222FCI limit with higher momentum corrections from ref. 174. -75.28127222FCI limit with higher momentum corrections from ref. 174. -100.001817333MRCI calculation from ref. 175.